# JoIN: Joint GANs Inversion for Intrinsic Image Decomposition

### v2

@@ 5.4 Ablation studies
In this section, we illustrate the benefits of each component of our method.
@@
Advantage of independently trained GANs. Training GANs independently makes our pipeline modular and flexible: different components can be trained with different data sources, and new components can be easily introduced in the forward model without re-training the entire pipeline. We showcase such use case by adding the specular component to the albedo-shading forward model (Fig. 6, 7, 3, 4).
@@
We also show that it is easier to train and invert GANs independently, further preventing cross-contamination during inversion. To demonstrate this, we train a single multi-component GAN that learns to generate both albedo and shading images simultaneously by producing a 6-channel output: 3 channels each for albedo and shading. We increase the number of trainable parameters by a factor of two to keep the capacity of the model comparable with our method.
@@
First, as shown in Fig. 10, such joint training leads to inferior image generation compared to independently-trained albedo and shading GANs. With a single GAN, albedo and shading are leaking into each other, this is also measured by higher FID scores.
Second, having the same network generating both albedo and shading means that the information is entangled up to the last layer, leading to cross-contamination during inversion. This can be seen in Fig. 11. While the image reconstruction quality is equivalent, the inversion is better using multiple GANs. Our observation is also supported with the quantitative results provided in Tab. 2.
Table 4: Effectiveness of kNN Loss in GAN Inversion
On CelebA-HQ test set Inversion on single GAN trained on FFHQ
LPIPSnormal-$\downarrow$ MSEnormal-$\downarrow$ PSNRnormal-$\uparrow$
Optimization [38] 0.3841 0.0950 16.9077
Opt. + InDomain loss [87, 71] 0.3839 0.0941 16.9901
Opt. + kNN loss (Ours) 0.3789 0.0919 17.1202
Figure 10: Single vs Separated GANs. Training separate GAN for each component results in better generation quality as opposed to training a single GAN for learning both components.
@@
Figure 11: Single vs Separated GANs Inversion. Inversion on a single GAN trained to generate both albedo and shading leads to cross-contamination issues (note the color leakage in shading and shadow effects in albedo). We mitigate this issue by training independent GANs for each component.
@@
Effectiveness of kNN loss. We evaluate the effectiveness of our newly proposed kNN-loss. For a fair evaluation, we first provide experiments for the more general single GAN inversion setup: we use a StyleGAN2 model trained on natural face images from FFHQ dataset [35], and invert CelebA-HQ testset [34] images using optimization-based inversion [38]. As shown in Tab. 4, using kNN loss as a regularization during optimization improves the performance over direct optimization and optimization with InDomain loss.
Full method ablations. Finally, we present a comprehensive set of ablations for our joint inversion of multiple GANs. We provide qualitative results (Fig. 12) and quantitative evaluations for intrinsic image decomposition on the Lumos Faces dataset in Tab. 2, second part.
@@
We first evaluate our method using only direct Optimization (Opt.). We then add the InDomain Loss discussed in Sec. 3.4 and then replace it with our kNN loss. As can be seen in Tab. 2 the kNN Loss allows better recovery of the intrinsic components by maintaining the latent codes within domain, it also outperforms the in-domain loss.
@@
We then add the encoder initialization from Sec. 4.1 to the pipeline and finally add the PTI-based finetuning, without, then with, the discriminator loss (D Loss) as discussed in Sec. 4.2. Our evaluation shows that the encoder initialization brings a clear improvement in both reconstruction quality and component estimations. The PTI without D Loss does improve the reconstruction of the input image but degrades the performance of component estimation, whereas, with the D Loss, the PTI still improves the image reconstruction quality but also provides the best results for the decomposition. A qualitative visualization of this difference is visible in Fig. 13.
Figure 12: Ablation study on Lumos synthetic face dataset. Our decomposition improves as we (from left to right) replace In-Domain loss with kNN loss; initialize with and encoder; use PTI scheme to finetune generators; and use local D loss during the PTI fine-tuning step.
@@
Figure 13: Importance of local D loss. Local D loss preserves the generator priors by preventing the local manifold from distortions. Above we show IID on a real face image without (top) and with (bottom) local discriminator loss.
6 Discussion and Future Work
@@
In this paper, we introduced JoIN a new method to solve inverse imaging problems using a bank of GANs. We show it is possible to retain GANs distribution priors while jointly inverting several of them in the case of Intrinsic Image Decomposition. Nonetheless, our method inherits some drawbacks of GAN inversion methods. First, it is slower than feed-forward methods with runtimes of a few seconds. Second, currently, training GANs on unaligned distributions remains a complex task and our method can be limited by the GAN training quality, hence it is more adapted to distributions that GANs can model well such as face images. In the future, as new generative models emerge, we believe our method will be applicable to broader datasets and could be tested on a wider variety of problems, possibly extending the ideas beyond images.
Supplementary Material
@@
We provide following as supplementary material to support the discussion and results of the main paper:
@@
  • In section 7, we discuss the intrinsic image decomposition experiments on primeshapes dataset. We come up with such fully synthetic dataset of simple shapes with variations of lights and materials in order to demonstrate the efficacy of our method with the two-component model (albedo and shading).
@@
  • In Sec. 8, we discuss our pipeline for generating and processing our datasets. Specifically, we discuss how we generate Primeshapes and Materials synthetic datasets using blender in python, and pre-processing steps involved for Lumos synthetic face dataset.
@@
  • In Sec. 10 we provide additional results for experiments discussed in the main paper to present supporting evidence for the claims made. In particular, we provide additional qualitative results for synthetic and real materials (Fig. 1617) and qualitative results for synthetic and real faces (Fig. 181921). Additional qualitative comparisons with competing methods on real faces are provided in Fig. 20. We also provide additional results for image relighting on faces using our method in Fig. 22.
7 Experiments on Primeshapes data
@@
[…]
Tone-mapping. Since the renderings generated using blender are in linear space, we apply suitable tone-mapping to each of the image components in order to use them in the GAN/encoder training. Our choice of tone-mapping functions is standard, and the same for both Primeshapes and Materials datasets. For albedo, we use gamma/sRGB tone-mapping with $\gamma=2.4$, while for shading and specular, we use Reinhard global tone-mapping [62] as these signals are inherently high-dynamic range.
8 Rendered Datasets details
Figure 15: Components available in our dataset. (Image source: Blender documentation [10]).
Data preparation pipeline. We use the blender python package [10] to generate both the Primeshapes and Materials synthetic datasets. We make use of rendering and passes functionalities in blender python to render $11$ components for each scene. Fig. 15 shows the relationship between these $11$ components, and how they can be combined to obtain the image. We keep the pre-processing steps the same for both the Primeshapes and Materials datasets: we use the diffuse color component as albedo; we add diffuse direct and diffuse indirect to obtain shading; and for specular, we multiply glossy color with the sum of glossy direct and glossy indirect. We ignore the transmission, emission, and environment components in our inversion as they do not contribute significantly to the final image in our setting. Thus our composed image, used as the target for inversion, is obtained with:
8.1 Primeshapes Details
@@ 8.2 Materials Details
9 Faces Dataset
@@
The Lumos dataset [79] provides thousands of synthetic face images captured in virtual lightroom created inside a 3D design software. For each image, various intrinsic sub-components such as albedo, shadow, sub-surface scattering, specular etc. are also provided. We combine these sub-components to obtain albedo, shading, and specular maps used to train our GANs. We leverage the facial image alignment method used for the FFHQ dataset [34, 35] to pre-process each image in the Lumos dataset. Such pre-processing ensures that all the facial images are aligned in the same fashion, thus improving GAN training and allowing us to test our method on the real images from FFHQ.
@@
Data preparation pipeline for faces. The Lumos synthetic dataset provides several image components for each face such as shadow, sub-surface scattering, coat, and sheen. While we use the albedo component directly from the dataset, we use the following strategy to obtain the shading and specular components for each image and train our GANs:
[…]
We continue to use the Eq. 14 for faces as well in order to combine albedo, shading, and specular to get the final image. Our tone-mapping functions for the Lumos dataset are similar to the Materials dataset: for albedo we use gamma/sRGB tone-mapping with $\gamma=2.4$, while for shading and specular, we use Reinhard tone-mapping [62].
10 Additional Results.
Qualitative Results on Faces. We provide additional results for decomposition into $2$ and $3$ components on synthetic face images in Fig. 18 and Fig. 19 respectively. As seen in the main paper, our model is able to recover albedo, shading, and specular components faithfully on synthetic images from the Lumos dataset. Further, we also provide additional comparisons with other methods in Fig. 20, along with the results on the real images from the FFHQ dataset in Fig. 21. As discussed in the main paper, our method is able to preserve the GAN priors and separate the image into plausible components, without copying shading into the albedo image.
@@
Image Relighting on Faces. As discussed in the main paper, our modular approach allows leveraging latent space editing capabilities of GANs to modify each image component separately, making it suitable for image modification such as relighting applications. We show additional image relighting results on real faces in Fig. 22. After obtaining the inversions for albedo, shading, and specular, we use an off-the-shelf latent space editing method SeFa [68] on the shading generator to modify the shading image while keeping other components constant. StyleGAN’s disentangled representation allows the identity of the person to remain relatively constant as we modify the shading image. As seen in Fig. 22, our approach can produce varied novel lighting.
Figure 22: Additional Face relighting results. Once the IID is obtained, we modify the shading component using GAN editing technique [68] on the shading generator obtaining plausible relighting edits.

### v3

@@ 5.4 Ablation Studies
Advantage of independently trained GANs. Training GANs independently makes our pipeline modular and flexible: different components can be trained with different data sources, and new components can be easily introduced in the forward model without re-training the entire pipeline. We showcase such use case by adding the specular component to the albedo-shading forward model (Fig. 6, 7(a), 5(a), 5(b)).
@@
We also show that it is easier to train and invert GANs independently, further preventing cross-contamination during inversion. To demonstrate this, we train a single multi-component GAN that learns to generate both albedo and shading images simultaneously by producing a 6-channel output: 3 channels each for albedo and shading. We increase the number of trainable parameters by $2\times$ for fair comparison with our method.
@@
First, as shown in Fig. 9(a), such joint training leads to inferior image generation compared to independently-trained albedo and shading GANs. With a single GAN, albedo and shading are leaking into each other, this is also measured by higher FID scores. Second, having the same network generating both albedo and shading means that the information is entangled up to the last layer, leading to cross-contamination during inversion as shown in Fig. 9(b). Our observation is also supported by the quantitative results provided in Tab. 4.
(a) Single GAN vs separated GANs as prior.
(b) GAN inversion on single vs separated GAN priors.
@@
Figure 9: (a) Training separate GAN for each component results in better generation quality as opposed to training a single GAN for learning both components. (b) Inversion on a single GAN trained to generate both albedo and shading leads to cross-contamination issues (note the color leakage in shading and shadow effects in albedo). We mitigate this issue by training independent GANs for each component.
Comparisons with DoubleDIP. We provide qualitative comparisons with doubleDIP (Gandelsman et al., 2019) in Fig. 10(a). As shown, untrained priors-based doubleDIP results in severe cross-contamination and loss of details on both the synthetic and real datasets as the underlying assumption of dissimilarity across the components does not hold. Our method produce accurate decompositions in both the cases.
@@
Effectiveness of kNN loss. To evaluate the effectiveness of our newly proposed kNN-loss, we first provide experiments for the more general single GAN inversion setup: we use a StyleGAN2 model trained on natural face images from FFHQ dataset (Karras et al., 2019), and invert CelebA-HQ testset (Karras et al., 2018) images using optimization-based inversion (Karras et al., 2020). As shown in Tab. 3, using kNN loss as a regularization improves the performance over direct optimization and optimization with in-domain loss. Further, we verify the effectiveness of kNN loss for IID task using joint inversion on synthetic faces. As shown in Fig. 10(b), kNN loss prevents cross contamination which is prevalent when using in-domain loss (notice glasses and shadows appearing in albedo), and improves the detail preservation (note mouth, beard, and hairstyle). Such improvements are supported by superior quantitative performance in Tab. 4.
(a) Comparisons with DoubleDIP.
(b) Effectiveness of kNN loss in joint inversion.
Figure 10: Ablation Results for (a) comparisons with DoubleDIP, and (b) effectiveness of kNN loss in joint inversion.
@@
Full method ablations. Finally, we present a comprehensive set of ablations for our joint inversion of multiple GANs. We provide qualitative results (Fig. 11(a)) and quantitative evaluations for intrinsic image decomposition on the Lumos Faces dataset in Tab. 4, second part. We first evaluate our method using only direct Optimization (Opt.). We then add the in-domain loss discussed in Sec. 3.4 and then replace it with our kNN loss. As can be seen in Tab. 4 the kNN Loss allows better recovery of the intrinsic components by maintaining the latent codes within domain, it also outperforms the in-domain loss.
@@
We then add the encoder initialization from Sec. 4.1 to the pipeline and finally add the generator fine-tuning (PTI), without, then with, the D Loss as discussed in Sec. 4.2. Our evaluation shows that the encoder initialization brings a clear improvement in both reconstruction quality and component estimations. The PTI without D Loss does improve the reconstruction of the input image but performs poorly in component estimation, whereas, with the D Loss, the PTI provides the best results for the decomposition while still improving the image reconstruction quality. A qualitative visualization of this difference is visible in Fig. 11(b).
(a) Full method ablations on synthetic faces.
(b) Ablation on importance of local D loss.
@@
Figure 11: (a) Our decomposition improves as we (from left to right) replace in-domain loss with kNN loss; initialize with and encoder; fine-tune generators using PTI; and use local D loss during the PTI fine-tuning step. (b) Local D loss preserves the generator priors by preventing the local manifold from distortions. Above we show IID on a real face image without (top) and with (bottom) local discriminator loss.
5.5 Storage, Memory, and Computational Costs
Runtime. Our approach adapts efficient GAN inversion techniques ($w$-optimization, encoder-based initialization, and Pivotal Tuning Inversion (PTI)) for joint GAN inversion, with our novel kNN and local discriminator losses adding minimal overhead ($\approx 15\%$ and $\approx 10\%$ to the optimization and PTI runtimes, respectively). This is achieved through GPU-accelerated faiss library (Johnson et al., 2019) for nearest neighbor search, which takes <2<2ms to compute nearest neighbors each optimization step (for $k=50$). On a single NVIDIA A40, JoIN takes $190$ seconds for inference on a single image ($84$s for optimization, $1$s for encoder, $105$s for PTI). While slower than feed-forward methods like InverseRenderNet++ ($2$s) and PieNet ($1$s), JoIN is considerably faster than optimization-based methods like NextFace ($320$s), and multiple priors methods such as DoubleDIP ($878$s), UIDNet ($950$s) and diffusion-based DMPlug ($635$s). Crucially, JoIN is fully amortizable, enabling concurrent processing of $8$ images on a single NVIDIA A40, reducing per-image inference time to $24$ seconds ($11$s for optimization, <1<1s for encoder, $13$s for PTI).
Storage and Memory. JoIN requires $\approx 800$MB for model storage, lower than PieNet (22002200MB), NextFace (13001300MB), and InverseRenderNet++ (12001200MB). GPU memory usage during inference for our method is 44GB, which is on par with other competing methods.
In summary, JoIN strikes a favorable balance between accuracy and efficiency. While feed-forward methods like InverseRenderNet++ and PieNet offer faster inference, they lack the flexibility to adapt to different forward models without retraining, a key advantage of our method. Compared to optimization-based methods like NextFace and prior-based methods like UIDNet, JoIN achieves significant runtime reductions, particularly considering the amortization.
6 Limitations and Future Work
@@
In this paper, we introduced JoIN – a new method to solve inverse problem of image decomposition using a bank of GANs as priors. We show that it is possible to retain GANs distribution priors while jointly inverting several GANs at once. Nonetheless, our method inherits some drawbacks of GAN inversion methods. First, it is slower than feed-forward methods with runtimes of a few seconds (detailed discussion is in appendix Sec. 5.5). Second, currently, training GANs on unaligned distributions remains a complex task and our method can be limited by the GAN training quality, hence it is more adapted to distributions that GANs can model well such as face images. Our framework of using bank of GANs as a prior is not limited to only the problem of IID, and can be extended to similar inverse problems in other domains of engineering such as audio source separation as a part of future work. An interesting avenue for future work is exploring the replacement of GAN priors with diffusion models, which have shown impressive results in various image generation and inverse problems, to investigate their potential for improving the accuracy and robustness of intrinsic decomposition. Particularly for larger and more diverse datasets, ability of diffusion models to capture complex data distributions may prove advantageous. As new techniques emerge, we believe our method will be applicable to broader datasets and could be tested on a wider variety of problems.
Appendix
@@
We provide following as appendix to support the discussion and results of the main paper:
@@
  • In Sec. A, we discuss the intrinsic image decomposition experiments on primeshapes dataset. We come up with such fully synthetic dataset of simple shapes with variations of lights and materials in order to demonstrate the efficacy of our method with the two-component model (albedo and shading). The qualitative results are shown in Fig. 12.
@@
  • In Sec. B, we discuss our pipeline for generating and processing our datasets. Specifically, we discuss how we generate Primeshapes and Materials synthetic datasets using blender in python, and pre-processing steps involved for Lumos synthetic face dataset. We also discuss the responsible usage of the human data.
  • In Sec. C, we discuss the responsible use of human data and other ethical considerations pertaining to our work.
@@
  • In Sec. D we provide additional results for experiments discussed in the main paper to present supporting evidence for the claims made. In particular, we provide:
    • Additional qualitative results for synthetic and real materials in Fig. 1415.
    • Additional qualitative results for synthetic and real faces in Fig. 161718.
    • Additional qualitative comparisons with competing methods on real faces in Fig. 19.
    • Additional results for image relighting on faces using our method in Fig. 20.
Appendix A Experiments on Primeshapes data
Appendix B Details of Rendered Datasets
Data preparation pipeline. We use the blender python package Blender Online Community (2023) to generate both the Primeshapes and Materials synthetic datasets. We make use of rendering and passes functionalities in blender python to render $11$ components for each scene. Fig. 13 shows the relationship between these $11$ components, and how they can be combined to obtain the image. We keep the pre-processing steps the same for both the Primeshapes and Materials datasets: we use the diffuse color component as albedo; we add diffuse direct and diffuse indirect to obtain shading; and for specular, we multiply glossy color with the sum of glossy direct and glossy indirect. We ignore the transmission, emission, and environment components in our inversion as they do not contribute significantly to the final image in our setting. Thus our composed image, used as the target for inversion, is obtained with:
final image $=\textnormal{shading}\cdot\textnormal{albedo}+\textnormal{specular}$ (13)
@@ Appendix B Details of Rendered Datasets
Inverse color-transform operation (tone-mapping). Since the renderings generated using blender are in linear space, we convert them in the color space by applying suitable inverse color-transform operation (more widely known as tone-mapping in the literature) to each of the image components in order to use them in the GAN/encoder training. Our choice of tone-mapping functions is standard, and the same for both Primeshapes and Materials datasets. For albedo, we use gamma/sRGB tone-mapping with $\gamma=2.4$, while for shading and specular, we use Reinhard global tone-mapping Reinhard et al. (2002) as these signals are inherently high-dynamic range.
B.1 Primeshapes Synthetic Dataset
@@ B.2 Materials Synthetic Dataset
Figure 13: Components available in our dataset. (Image source: Blender documentation Blender Online Community (2023)).
B.3 Lumos Synthetic Faces Dataset
@@
The Lumos dataset (Yeh et al., 2022) provides thousands of synthetic face images captured in virtual lightroom created inside a 3D design software. For each image, various intrinsic sub-components such as albedo, shadow, sub-surface scattering, specular etc. are also provided. We combine these sub-components to obtain albedo, shading, and specular maps used to train our GANs. We leverage the facial image alignment method used for the FFHQ dataset (Karras et al., 2018; 2019) to pre-process each image in the Lumos dataset. Such pre-processing ensures that all the facial images are aligned in the same fashion, thus improving GAN training and allowing us to test our method on the real images from FFHQ.
The Lumos synthetic dataset provides several image components for each face such as shadow, sub-surface scattering, coat, and sheen. While we use the albedo component directly from the dataset, we use the following strategy to obtain the shading and specular components for each image and train our GANs:
specular $=\textnormal{specular}+\textnormal{sheen}+\textnormal{coat}$ (15)
We continue to use the Eq. 13 for faces as well in order to combine albedo, shading, and specular to get the final image. Our tone-mapping functions for the Lumos dataset are similar to the Materials dataset: for albedo we use gamma/sRGB tone-mapping with $\gamma=2.4$, while for shading and specular, we use Reinhard tone-mapping (Reinhard et al., 2002).
Appendix C Broader Impact Concerns
C.1 Responsible Use of Human Data
We perform our experiments on images from four different domains: primeshapes, materials, faces, and indoor scenes. For each case, we use fully synthetic datasets for training the GAN models on individual intrinsic components, and use synthetic and/or real data for testing and evaluation of our model. None of the synthetic datasets we use contain any human subject data or personally identifiable information. The only scenario where we potentially use human subject data is for evaluation of our method on real faces using FFHQ dataset.
In particular, for faces, we use Lumos synthetic face dataset for training, and a small number of real images from FFHQ dataset Karras et al. (2018; 2019) for evaluating our model. FFHQ dataset contains $70,000$ images of human faces collected from an online image sharing platform named Flickr. This dataset is diverse, featuring people of various ages, ethnicity, and backgrounds. It even includes a wide range of accessories like glasses, sunglasses, and hats. The source for these images was Flickr, so it’s important to be aware of potential biases present on that platform that maybe inherited by the dataset. During the collection of the dataset, it was made sure that only those images with permissive licenses are included. Finally, to ensure the dataset’s quality, automated filters and human reviewers (via Amazon Mechanical Turk) removed any irrelevant images like statues or pictures of pictures. It is important to note that this dataset is not intended to be used for any facial recognition technology. Licensing and usage information about the dataset is available on its webpage [^1] https://github.com/NVlabs/ffhq-dataset. We are committed to the responsible usage of human subjects data, and made sure we follow all the licensing requirements and policies in our experiments with FFHQ dataset.
C.2 Ethical Considerations
Although our method focuses on intrinsic image decomposition, as discussed in Sec. 5, one can potentially combine our approach with GAN’s latent space image editing to achieve image relighting on human faces. While such relighting does not change the underlying appearance of the face, potential broader impact concerns still exist. Realistic relighting could be misused to misrepresent the context of an image by subtly altering the perceived environment, potentially misleading viewers. Furthermore, even without altering facial features, our method might inadvertently reinforce biases present in the training data regarding lighting conditions and their association with certain demographics. Finally, relighting could complicate the authentication of images, making it more challenging to detect subtle forms of manipulation. Therefore, we advocate for responsible development and deployment of relighting technology, alongside ongoing research into detection methods, to mitigate these risks and maintain trust in visual media. Transparency about our method’s limitations is crucial for fostering informed discussion and responsible use.
Appendix D Additional Results
Qualitative results on faces. We provide additional results for decomposition into $2$ and $3$ components on synthetic face images in Fig. 16 and Fig. 17 respectively. As seen in the main paper, our model is able to recover albedo, shading, and specular components faithfully on synthetic images from the Lumos dataset. Further, we also provide additional results on the real images from the FFHQ dataset in Fig. 18, along with the comparisons with other methods in Fig. 19. As discussed in the main paper, our method is able to preserve the GAN priors and separate the image into plausible components, without copying shading into the albedo image.
Image relighting on faces. As discussed in the main paper, our modular approach allows leveraging latent space editing capabilities of GANs to modify each image component separately, making it suitable for image modification such as relighting applications. We show additional image relighting results on real faces in Fig. 20. After obtaining the inversions for albedo, shading, and specular, we use an off-the-shelf latent space editing method SeFa (Shen & Zhou, 2021) on the shading generator to modify the shading image while keeping the other two components constant. StyleGAN’s disentangled representation allows the identity of the person to remain relatively constant as we modify the shading image. As seen in Fig. 20, our approach can produce varied novel lighting.
@@ Appendix D Additional Results
Figure 20: Additional Face relighting results. Once the IID is obtained, we modify the shading component using GAN editing technique Shen & Zhou (2021) on the shading generator obtaining plausible relighting edits.